# Provable Benefit of Curriculum in Transformer Tree-Reasoning Post-Training

### v1

@@ 1 Introduction
To bridge this gap, we observe that in post-training, “performance” is naturally measured by the probability of generating the correct CoT, while “difficulty” is often quantified by the pass-rate, i.e., the probability that the base model already produces the correct CoT (tong2024dartmathdifficultyawarerejectiontuning; parashar2025curriculum). A principled formalization of this “difficulty” uses the coverage coefficient foster2025goodfoundationnecessaryefficient $\|\tfrac{\pi^{\star}}{\pi_{\text{ref}}}\|_{\infty}:=\sup_{x}\|\tfrac{d\pi^{\star}(\cdot\mid x)}{d\pi_{\text{ref}}(\cdot\mid x)}\|_{L^{\infty}(\pi_{\text{ref}}(\cdot\mid x))}<\infty$ (assuming absolute continuity), which bounds how much rarer a correct CoT can be under the base model’s predictive policy $\pi_{\text{ref}}$ compared to the correct CoT policy $\pi^{\star}$, and thus inversely control the pass-rate. In this view, a curriculum corresponds to intermediate policies $\pi_{0}^{\star}=\pi_{\mathrm{ref}},\pi_{1}^{\star},\ldots,\pi_{K}^{\star}=\pi^{\star}$, with the “difficulty” $\|\tfrac{\pi_{k}^{\star}}{\pi_{\text{ref}}}\|_{\infty}$ nondecreasing in $k\in[K]$.
@@ 2 Theoretical Framework for Curriculum Post-training on reasoning trees
Preliminaries and Notations. For each prompt $x$, policies are conditional probability measures $\pi(\cdot\mid x)$ on the output space. Whenever $\pi_{k^{\prime}}(\cdot\mid x)\ll\pi_{k}(\cdot\mid x)$, let $r^{x}_{k^{\prime},k}(o):=\tfrac{d\pi_{k^{\prime}}(\cdot\mid x)}{d\pi_{k}(\cdot\mid x)}(o)$. We write $\|\frac{\pi_{k^{\prime}}}{\pi_{k}}\|_{\infty}:=\sup_{x}\|r^{x}_{k^{\prime},k}\|_{L^{\infty}(\pi_{k}(\cdot\mid x))}=\sup_{x}\operatorname*{esssup}_{o\sim\pi_{k}(\cdot\mid x)}r^{x}_{k^{\prime},k}(o)$. For a function $f$, $\|f\|_{\infty}:=\operatorname*{esssup}_{z}|f(z)|$. Landau symbols follow standard conventions. For $\mathbf{a}\in\mathbb{R}^{m}$, $\operatorname{softmax}(\mathbf{a})_{i}:=\exp(a_{i})/\sum_{j}\exp(a_{j})$, and with temperature $\beta>0$ we write $\operatorname{softmax}(\mathbf{a}/\beta)$. Finally, $\operatorname{ReLU}(t):=\max\{0,t\}$ elementwise.
@@
Building on the introduction’s discussion of the pass-rate as a difficulty measure in post-training reasoning (tong2024dartmathdifficultyawarerejectiontuning; parashar2025curriculum), where $\|\tfrac{\pi^{\star}}{\pi_{\mathrm{ref}}}\|_{\infty}$ serves as a natural proxy for difficulty, we state the following theorem.
@@
Consider a curriculum of $K$ tasks with base model $\pi_{\mathrm{ref}}:=\pi^{\star}_{0}$. For $0\leq k<k^{\prime}\leq K$, let $C_{k^{\prime},k}:=\|\frac{\pi_{k^{\prime}}^{\star}}{\pi_{k}^{\star}}\|_{\infty}$. Assume absolute continuity and finiteness $C_{k^{\prime},k}<\infty$ for all $k<k^{\prime}$. Suppose the optimal policies $\{\pi^{\star}_{k}\}_{k=0}^{K}$ satisfy the stepwise range and a multiplicative alignment for all $k<k;$
$C_{k+1,k}=\Theta(C^{\star}),$ (1)
$C_{k^{\prime},k}=\Theta\!\Big(\prod_{r=k}^{k^{\prime}-1}C_{r+1,r}\Big),$ (2)
for some positive constant $C^{\star}>1$. Further assume a complexity–mismatch alignment
$\mathcal{C}(\pi^{\star}_{k^{\prime}}\mid\pi^{\star}_{k})=\Theta\!\big(C_{k^{\prime},k}\big),$ (3)
up to harmless logarithmic factors in a confidence parameter. Then:
  • Direct estimation of $\pi^{\star}_{K}$ from $\pi^{\star}_{0}$ incurs $\mathcal{C}_{\mathrm{direct}}:=\mathcal{C}(\pi^{\star}_{K}\mid\pi^{\star}_{0})=\Theta\!\big(C_{K,0}\big)=\Theta\!\big((C^{\star})^{K}\big)$.
  • Stepwise curriculum learning incurs total complexity $\mathcal{C}_{\mathrm{curriculum}}:=\sum_{k=0}^{K-1}\mathcal{C}(\pi^{\star}_{k+1}\mid\pi^{\star}_{k})=\Theta\!\big(\sum_{k=0}^{K-1}C_{k+1,k}\big)=\Theta(KC^{\star})$.
@@
Related Theoretical Context. Indeed, the assumptions in parashar2025curriculum closely mirror those above in the case of Approximate Policy Iteration settings. We discuss this comparison in App. B. Furthermore, for linearly realizable Markov Decision Processes, leveraging the spanner-sampling framework of foster2025goodfoundationnecessaryefficient, we can also prove an exponential improvement in inference-time computational complexity (Thm. thm:curriculum-spanner) by similar assumptions; see App. B for details.
@@ 2.1 Transformer Reasons over a States-Conditioned Reasoning Tree
A detailed version of Def. 1 (Def. 3) appears in App. D. To rule out ambiguities where different index paths represent the same task, we adopt the uniqueness assumption below.
@@
Transformer ($\operatorname{TF}(\cdot;\mathbf{W})$) as Learning Model. Let $\mathbf{U}=[\bm{\mu}_{1},\ldots,\bm{\mu}_{K},\bm{\mu}_{\mathrm{EOS}}]\in\mathbb{R}^{d_{\mathrm{X}}\times(K+1)}$ be the embedding vocabulary, where the $k$-th column is the embedding of token $k$, and $\bm{\mu}_{\mathrm{EOS}}$ the embedding of $\mathrm{EOS}$. Following li2025iclcot, positional encodings $\mathbf{p}_{1},\ldots,\mathbf{p}_{d+L+1}$ are mutually orthogonal and satisfy $\mathbf{p}_{i}\perp\mathbf{U}$. They are concatenated with token embeddings as $\mathbf{E}[x_{m}]=[{\bm{\mu}^{x_{m}}}^{\top},{\mathbf{p}_{m}}^{\top}]^{\top}$, $\mathbf{E}[\mathrm{EOS}]=[{\bm{\mu}_{\mathrm{EOS}}}^{\top},{\mathbf{p}_{d+1}}^{\top}]^{\top}$, and $\mathbf{E}[z_{l}]=[{\bm{\mu}^{z_{l}}}^{\top},{\mathbf{p}_{l+d+1}}^{\top}]^{\top}$, $\forall m\in[d],l\in[L+1]$. The transformer $\operatorname{TF}(\cdot;\mathbf{W})$ aims to recover the functions defined in the tree graph $f\in\mathcal{F}_{G}$, with embedded input $\mathbf{E}[x_{1}],\ldots,\mathbf{E}[x_{d}],\mathbf{E}[\mathrm{EOS}]\in\mathbb{R}^{d_{\mathrm{E}}}$, where $d_{\mathrm{E}}=2d_{\mathrm{X}}$. Denote $\mathbf{E}[z_{0}]:=\mathbf{E}[\mathrm{EOS}]$, the transformer performs next-token prediction via:
$\operatorname{TF}(\mathbf{E}[x_{1}],\ldots,\mathbf{E}[x_{d}],\mathbf{E}[z_{0}],\mathbf{E}[z_{1}],\ldots,\mathbf{E}[z_{l-1}];\mathbf{W})=\mathbf{E}[z_{l}]$
@@
  • $\mathbf{R^{\mathcal{F}_{S_{\star}}}_{\mathbf{x}}(\cdot)}$: for the curriculum subtask family $\mathcal{F}_{S_{\star}}=\{f_{S_{\star}^{1}},\ldots,f_{S_{\star}^{k^{\star}}}\}$ per Def. 1, returns $1$ iff the pre-$\mathrm{EOS}$ token equals $\bm{\mu}^{f_{S_{\star}^{\ell}}(\mathbf{x})}$ for some $\ell\in[k^{\star}]$; else 0.
@@ 3 Conclusion, Discussion, and Future Work
A limitation is that our 2S-ART (PART) abstraction is restrictive. For instance, datasets such as Blocksworld involve $\Phi_{l}(x_{i_{l}^{1}},x_{i_{l}^{2}})$ (choosing two states and stacking one on the other), beyond our one-state case. We believe the theory extends with more complex treatment. While our analysis focused on transformer post-training, the benefit of curricula in avoiding exponential bottlenecks is model-agnostic, as highlighted in Thm. 1 and Sec. B. Empirical studies also report easy-to-hard curricula to be effective in diffusion models (kim2025orderMD; kim2025curriculumDM; yu2023whiteBoxTransformers), and it is an interesting direction to investigate whether their underlying principles align with our theorem.
@@ Appendix A Additional Related Work
Beyond expressiveness, recent studies examine how transformers trained with CoT evolve during optimization. Early work explored emergent in-context learning and convergence under multi-head attention (hahn2023theory; huang2023context; yang2024incontextlearningrepresentationscontextual; zhang2024trained), while subsequent analyses provide provable characterizations of induction heads and sparse token mechanisms (chen2024unveilinginductionheadsprovable). Of particular relevance are studies on CoT-specific training dynamics, which analyze nonlinear or single-layer transformers solving structured tasks such as parity (wen2024sparse; kim2025parity; li2025iclcot). More recent work extends these insights to multi-step optimization and regular language recognition, uncovering implicit biases and algorithmic behaviors in gradient descent (huang2025transformers; huang2025how; huang2025CoT). wen2024sparsedependencesparseattention; kim2025parity investigated the learning of transformers on the parity task with CoT, building on which yin2025datashiftshurtcot studied the impact of data shift. yang2025symbolic further demonstrated that transformers can implement both forward and backward tree-structured symbolic reasoning using two attention heads. However, these works abstract away real-world sampling process during inference, relying instead on atypical deterministic prediction of internal tokens/sentences, and they primarily analyze pre-training rather than post-training.
Appendix B Cases beyond transformer gradient dynamics over Tree
@@
Appendix C Auxiliary Lemmas
@@
Appendix D Details and Proofs of Autoregressive Reasoning Tree

### v2

@@ 1 Introduction
To bridge this gap, we observe that in post-training, “performance” is naturally measured by the probability of generating the correct CoT, while “difficulty” is often quantified by the pass-rate, i.e., the probability that the base model already produces the correct CoT (tong2024dartmathdifficultyawarerejectiontuning; parashar2025curriculum). A principled formalization of this “difficulty” uses the coverage coefficient foster2025goodfoundationnecessaryefficient $\|\tfrac{\pi^{\star}}{\pi_{\text{ref}}}\|_{\infty}:=\sup_{x}\|\tfrac{d\pi^{\star}(\cdot\mid x)}{d\pi_{\text{ref}}(\cdot\mid x)}\|_{L^{\infty}(\pi_{\text{ref}}(\cdot\mid x))}<\infty$111i.e., the $\ell_{\infty}$-norm of the Radon–Nikodym derivative of $\pi^{\star}$ w.r.t. $\pi_{\text{ref}}$. If $\mathcal{X}\times\mathcal{Y}$ are discrete space, it holds that $\|\tfrac{\pi^{\star}}{\pi_{\text{ref}}}\|_{\infty}=\sup_{x\in\mathcal{X},y\in\mathcal{Y}}\frac{\pi^{\star}(y|x)}{\pi_{\text{ref}}(y|x)},$ (assuming absolute continuity), which bounds how much rarer a correct CoT can be under the base model’s predictive policy $\pi_{\text{ref}}$ compared to the correct CoT policy $\pi^{\star}$, and thus inversely control the pass-rate. In this view, a curriculum corresponds to intermediate policies $\pi_{0}^{\star}=\pi_{\mathrm{ref}},\pi_{1}^{\star},\ldots,\pi_{K}^{\star}=\pi^{\star}$, with the “difficulty” $\|\tfrac{\pi_{k}^{\star}}{\pi_{\text{ref}}}\|_{\infty}$ nondecreasing in $k\in[K]$.
@@ 2 Theoretical Framework for Curriculum Post-training on reasoning trees
Preliminaries and Notations. For each prompt $x$, a policy $\pi$ is a conditional probability measure $\pi(\cdot\mid x)$ over the output space $\mathcal{O}$, where $o\in\mathcal{O}$ denotes a CoT trajectory. For two policies $\pi$ and $\pi^{\prime}$, if $\pi^{\prime}(\cdot\mid x)\ll\pi(\cdot\mid x)$ for all $x$, we write $\bigl\|\tfrac{\pi^{\prime}}{\pi}\bigr\|_{\infty}:=\sup_{x\in\mathcal{X}}\bigl\|\tfrac{d\pi^{\prime}(\cdot\mid x)}{d\pi(\cdot\mid x)}\bigr\|_{L^{\infty}(\pi(\cdot\mid x))}=\sup_{x\in\mathcal{X}}\operatorname*{ess\,sup}_{o\sim\pi(\cdot\mid x)}\tfrac{d\pi^{\prime}(\cdot\mid x)}{d\pi(\cdot\mid x)}(o)$. Landau symbols follow standard conventions. For $\mathbf{a}\in\mathbb{R}^{m}$, $\operatorname{softmax}(\mathbf{a})_{i}:=\exp(a_{i})/\sum_{j}\exp(a_{j})$, and with temperature $\beta>0$ we write $\operatorname{softmax}(\mathbf{a}/\beta)$. Finally, $\operatorname{ReLU}(t):=\max\{0,t\}$ elementwise.
@@
As discussed in the introduction, pass rate is the operative notion of difficulty in post-training reasoning (tong2024dartmathdifficultyawarerejectiontuning; parashar2025curriculum), and the inverse of $\|\tfrac{\pi^{\star}}{\pi_{\mathrm{ref}}}\|_{\infty}$ lower-bounds the pass rate and thus serves as a natural proxy for difficulty to master the target task. We formalize the discussions in the introduction as the theorem below.
@@
Consider a curriculum of $K$ tasks $\pi_{0}^{\star},\pi_{1}^{\star},\ldots,\pi_{K}^{\star}$ with base model $\pi_{\mathrm{ref}}:=\pi^{\star}_{0}$ and target task $\pi^{\star}:=\pi_{K}^{\star}$. Denote the learning complexity [^2] The concrete complexity is left to be defined for certain specific scenario. For instance, it can be the sample Complexity of finetuning algorithms (Thm. 3), or the reward-oracle query complexity of test-time scaling (Thm. 4). from $\pi_{k}^{\star}$ to $\pi_{k^{\prime}}^{\star}$ as $\mathcal{C}(\pi^{\star}_{k^{\prime}}\mid\pi^{\star}_{k})$. For any $0\leq k<k^{\prime}\leq K$, suppose $\|\frac{\pi_{k^{\prime}}^{\star}}{\pi_{k}^{\star}}\|_{\infty}<\infty$ and
$\|\frac{\pi_{k+1}^{\star}}{\pi_{k}^{\star}}\|_{\infty}=\Theta(C^{\star}),$ (1)
$\|\frac{\pi_{k^{\prime}}^{\star}}{\pi_{k}^{\star}}\|_{\infty}=\Theta\!\Big(\prod_{r=k}^{k^{\prime}-1}\|\frac{\pi_{r+1}^{\star}}{\pi_{r}^{\star}}\|_{\infty}\Big),$ (2)
$\mathcal{C}(\pi^{\star}_{k^{\prime}}\mid\pi^{\star}_{k})=\widetilde{\Theta}\!\big(\|\frac{\pi_{k^{\prime}}^{\star}}{\pi_{k}^{\star}}\|_{\infty}\big),$ (3)
[…]
  • Direct estimation of $\pi^{\star}_{K}$ from $\pi^{\star}_{0}$ costs $\mathcal{C}_{\mathrm{direct}}:=\sum_{k=0}^{K-1}\mathcal{C}(\pi^{\star}_{k+1}\mid\pi^{\star}_{k})=\Theta\!\big((C^{\star})^{K}\big)$.
  • Stepwise curriculum costs $\mathcal{C}_{\mathrm{curriculum}}:=\sum_{k=0}^{K-1}\mathcal{C}(\pi^{\star}_{k+1}\mid\pi^{\star}_{k})=\Theta(KC^{\star})$.
@@
Related Theoretical Context. Indeed, the assumptions in parashar2025curriculum closely mirror those above in the case of Approximate Policy Iteration settings. We discuss this comparison in App. E. Furthermore, for linearly realizable Markov Decision Processes, leveraging the spanner-sampling framework of foster2025goodfoundationnecessaryefficient, we can also prove an exponential improvement in inference-time computational complexity (Thm. 5) by similar assumptions; see App. E for details.
@@ 2.1 Transformer Reasons over a States-Conditioned Reasoning Tree
A detailed version of Def. 1 (Def. 3) appears in App. G. To rule out ambiguities where different index paths represent the same task, we adopt the uniqueness assumption below.
@@
Transformer ($\operatorname{TF}(\cdot;\mathbf{W})$) as Learning Model. Let $\mathbf{U}=[\bm{\mu}_{1},\ldots,\bm{\mu}_{K},\bm{\mu}_{\mathrm{EOS}}]\in\mathbb{R}^{d_{\mathrm{X}}\times(K+1)}$ be the embedding vocabulary, where the $k$-th column is the embedding of token $k$, and $\bm{\mu}_{\mathrm{EOS}}$ the embedding of $\mathrm{EOS}$. Following li2025iclcot, positional encodings $\mathbf{P}:=[\mathbf{p}_{1},\ldots,\mathbf{p}_{d+1}]$ are mutually orthogonal and satisfy $\mathbf{p}_{i}\perp\mathbf{U}$. They are concatenated with token embeddings as $\mathbf{E}[x_{m}]=[{\bm{\mu}^{x_{m}}}^{\top},{\mathbf{p}_{m}}^{\top}]^{\top}$, $\mathbf{E}[\mathrm{EOS}]=[{\bm{\mu}_{\mathrm{EOS}}}^{\top},{\mathbf{p}_{d+1}}^{\top}]^{\top}$, and $\mathbf{E}[z_{l}]=[{\bm{\mu}^{z_{l}}}^{\top},{\mathbf{p}_{i_{l}}}^{\top}]^{\top}$, $\forall m\in[d],l\in[L+1],i_{l}\in[d+1]$. The transformer $\operatorname{TF}(\cdot;\mathbf{W})$ aims to recover the functions defined in the tree graph $f\in\mathcal{F}_{G}$, with embedded input $\mathbf{E}[x_{1}],\ldots,\mathbf{E}[x_{d}],\mathbf{E}[\mathrm{EOS}]\in\mathbb{R}^{d_{\mathrm{E}}}$. Denote $\mathbf{E}[z_{-d}]:=\mathbf{E}[x_{1}],...,\mathbf{E}[z_{-1}]:=\mathbf{E}[x_{d}],\mathbf{E}[z_{0}]:=\mathbf{E}[\mathrm{EOS}]$, the transformer performs next-token prediction via:
$\operatorname{TF}(\mathbf{E}[z_{-d}],\ldots,\mathbf{E}[z_{-1}],\mathbf{E}[z_{0}],\mathbf{E}[z_{1}],\ldots,\mathbf{E}[z_{l-1}];\mathbf{W})=\mathbf{E}[z_{l}]$
@@
  • $\mathbf{R^{\mathcal{F}_{S_{\star}}}_{\mathbf{x}}(\cdot,\ell)}$: for the curriculum subtask family $\mathcal{F}_{S_{\star}}=\{f_{S_{\star}^{1}},\ldots,f_{S_{\star}^{k^{\star}}}\}$ per Def. 1, returns $1$ iff the pre-$\mathrm{EOS}$ token equals $\bm{\mu}^{f_{S_{\star}^{\ell}}(\mathbf{x})}$ for $\ell\in[k^{\star}]$; else 0.
@@ 2.1.2 Outcome Signal-based RL Finetunings by Gradient Descent
$\mathbf{K}^{\top}\mathbf{Q}\,=\,\begin{pmatrix}\bm{0}_{d_{\mathrm{X}}\times d_{\mathrm{X}}}&\bm{0}_{d_{\mathrm{X}}\times(d_{\mathrm{E}}-d_{\mathrm{X}})}\\ \bm{0}_{(d_{\mathrm{E}}-d_{\mathrm{X}})\times d_{\mathrm{X}}}&\mathbf{W}\end{pmatrix},\qquad\mathbf{V}\,=\,\begin{pmatrix}\bm{0}&\mathbf{I}_{d_{\mathrm{E}}-d_{\mathrm{X}}}\end{pmatrix}$, (10)
@@ 3 Conclusion, Discussion, and Future Work
A limitation is that our 2S-ART (PART) abstraction is restrictive. For instance, datasets such as Blocksworld involve $\Phi_{l}(x_{i_{l}^{1}},x_{i_{l}^{2}})$ (choosing two states and stacking one on the other), beyond our one-state case. We believe the theory extends with more complex treatment. While our analysis focused on transformer post-training, the benefit of curricula in avoiding exponential bottlenecks is model-agnostic, as highlighted in Thm. 1 and Sec. E. Empirical studies also report easy-to-hard curricula to be effective in diffusion models (kim2025orderMD; kim2025curriculumDM; yu2023whiteBoxTransformers), and it is an interesting direction to investigate whether their underlying principles align with our theorem.
@@ Appendix A Additional Related Work
Appendix B LLM Usage
In accordance with the conference policy on large language models (LLMs), we disclose that an LLM was used solely as a writing aid to improve the clarity and readability of certain paragraphs. The LLM did not contribute to research ideation, technical development, experimental design, analysis, or interpretation of results. All scientific ideas, methods, and conclusions presented in this paper are entirely the work of the authors.
Appendix C Ethics Statement
This work is purely theoretical and algorithmic in nature. It does not involve human subjects, personally identifiable data, or sensitive datasets, and therefore does not pose direct risks to privacy, security, or safety. Our study is conducted under the principles of academic integrity and transparency, adhering to the ICLR Code of Ethics. We further note that our methods and results do not introduce discriminatory or harmful applications, nor are they tied to any conflicts of interest or external sponsorship that may bias the research.
Appendix D Reproducibility Statement
We have taken concrete steps to ensure reproducibility of our results. All theoretical claims are accompanied by formal assumptions and complete proofs in the appendix. The complexity analyses and derivations are presented in full detail, ensuring transparency and verifiability. For experimental validations (if applicable), datasets are publicly available and preprocessing steps are explicitly described in the supplementary materials. To further support reproducibility, we provide pseudo-code and detailed algorithmic specifications in Alg. 231546, and will release anonymous source code in the supplementary materials to facilitate independent verification.
@@ Appendix D Reproducibility Statement
Beyond expressiveness, recent studies examine how transformers trained with CoT evolve during optimization. Early work explored emergent in-context learning and convergence under multi-head attention (hahn2023theory; huang2023incontext; yang2024incontextlearningrepresentationscontextual; zhang2024trained), while subsequent analyses provide provable characterizations of induction heads and sparse token mechanisms (chen2024unveilinginductionheadsprovable). Of particular relevance are studies on CoT-specific training dynamics, which analyze nonlinear or single-layer transformers solving structured tasks such as parity (wen2024sparse; kim2025parity; li2025iclcot). More recent work extends these insights to multi-step optimization and regular language recognition, uncovering implicit biases and algorithmic behaviors in gradient descent (huang2025transformers; huang2025how; huang2025CoT). wen2024sparsedependencesparseattention; kim2025parity investigated the learning of transformers on the parity task with CoT, building on which yin2025datashiftshurtcot studied the impact of data shift. yang2025symbolic further demonstrated that transformers can implement both forward and backward tree-structured symbolic reasoning using two attention heads. However, these works abstract away real-world sampling process during inference, relying instead on atypical deterministic prediction of internal tokens/sentences, and they primarily analyze pre-training rather than post-training.
Appendix E Cases beyond transformer gradient dynamics over Tree
@@
Appendix F Auxiliary Lemmas
@@
Appendix G Details and Proofs of Autoregressive Reasoning Tree